You are an equity analyst at Kowloon Bay Equities and are following the stock of a screwdriver manufacturer called TST Tools Inc.

On October 31st, 2019, TST Tools Inc. has 5,000 shares outstanding and based on your analysis of its performance, you expect TST Tools Inc. to report an earnings per share of $4.00 for the quarter ended on December 31st, 2019.

On November 30th, 2019 TST Tools Inc. files an ‘8-K’ announcement, in which it says that it has to take a one-time impairment charge for its Goodwill in Mong Kok Inc. of $8,000.

On December 31st, 2019, TST Tools reports an earnings per share of $2.50 for the quarter of October 1st, 2019 to December 31st, 2019.

Would you change your expectation of TST Tools. Inc.’s EPS after November 30th, 2019 and by how much? (2 points)

Did TST Tools meet your expectations on December 31st, 2019? In one sentence, please explain why or why not? (2 points)

Why did TST Tools Inc. record a big Goodwill charge of $8,000 on November 30th, 2019? Please explain in less than one sentence (3 points)
Answer #1

1. Net Income = 5000 *4 = 20,000

Net income after change = 20,000 - 8000 = 12000

EPS after change = 12000 /4000 = $3

Change = $4-$3 = $1

2. No ,it did not meet the expectations since the EPS expected post write off is $3 whereas the actual reporting is $2.5.

3. Goodwill and other intangible assets must be tested for impairment that the end of each reporting period to ascertain if the value for the asset still exists or there is a decline in the value.